# On the Isbell Problem

### v1

@@ 3. $(iii)$
Let $x\subseteq\kappa$ be an infinite subset having order type $\omega$ and $f:x\to 2$. Since $\mathbb{Q}_{\kappa}$ is Cohen preserving, by Lemma 4.22 we find $\vec{\nu}\in\mathcal{C}_{\kappa}\cap V$ such that $\bigcap_{\alpha\in{x}}{B}^{\alpha}_{{f}(\alpha)}\subseteq{X}_{\vec{\nu}_{0}}$. By the last remark of the previous paragraph, ${\varphi}^{-1}_{\alpha_{\gamma}}[{X}_{\vec{\nu}_{0}}]\in{\mathcal{U}}_{\alpha_{\gamma}}^{*}$, which implies ${\varphi}^{-1}_{\alpha_{\gamma}}[\bigcap_{\alpha\in{x}}B^{\alpha}_{{f}(\alpha)}]\in{\mathcal{U}}_{\alpha_{\gamma}}^{*}$. Thus, $\bigcap_{\alpha\in{x}}{D}^{\alpha}_{{f}(\alpha)}\in{\mathcal{U}}_{\alpha_{\gamma}}^{*}$. Note that this implies that any ultrafilter extending $\mathcal{U}_{\alpha_{\gamma}}$ is Tukey top. Since $\mathcal{U}_{\alpha_{\gamma}}=\mathcal{F}\cap V[H_{\alpha_{\gamma}}]\subseteq\mathcal{F}\subseteq\mathcal{U}$, we have that $\mathcal{U}$ is Tukey top.
@@ 8. Squares and products
In this section we present a different approach based on the existence of square sequences. Note that, working in the Sacks mode l, the quotient forcing $[\mathbb{Q}_{\kappa}:\mathbb{S}_{\kappa}]$ is naturally equivalent to a subposet of the countable support product $\bigotimes_{\beta\in S}\dot{\mathbb{P}}_{\kappa}(\dot{\mathcal{U}}_{\beta})$. In this section we provide a construction of a reduced product which keeps all the relevant properties of $\mathbb{Q}_{\kappa}$, although it looks quite different in what the support of the conditions refers. We focus only on the construction of such poset and prove that it has all the important properties required for the proof of the main theorem. The construction of the forcing is based on the next theorem which is interesting by itself.
@@
Suppose that $n\in\omega$ is a natural number and $\Box_{m}$ holds for all $m<n$ and $\mathtt{cof}(meager)=\aleph_{1}$ holds. Let $\langle\mathbb{P}_{\alpha}\colon\alpha\in\omega_{n}\rangle$ be a sequence of strongly proper, Cohen preserving posets of cardinality $\aleph_{1}$. Then there is a poset $\mathbf{P}$ which:
@@
If $V=L$ and $\vec{B}$ is an $\omega_{1}$-sequence of ordinals, then $L[\vec{B}]$ is a model of $\mathsf{ZFC}+\mathsf{GCH}$.
@@
The model $M_{a}$ is $L[K][S\cap\mathcal{P}(\mathrm{cl}(a))][\langle\langle\alpha,\mathbb{P}_{\alpha}\rangle\colon\alpha\in a\rangle]$. One point here is that this is a model of ZFC, since the set $S\cap\mathcal{P}(\mathrm{cl}(a))$ is injectively mapped into $Ord^{n+1}$ by the function $s$ and thereby canonically well-ordered. Items (1, 2) are obvious. For item (3), note that $S\cap\mathcal{P}(cl(a))$ is a countable set, so $M_{a}=L[\vec{B}_{a}]$ for some $\omega_{1}$-sequence $\vec{B}_{a}$ of ordinals and it computes $\aleph_{1}$ correctly by the choice of the sequence $K$. It follows from 8.10 that $M_{a}$ is a model of $\mathsf{CH}$. In the model $M_{a}$, let $\prec_{a}$ be the first well-ordering of the reals of ordertype $\omega_{1}$ in the natural constructibility well-order. For the last item, note that if $b\subseteq a$ are sets in $S$, then $M_{a}$ can compute $\mathrm{cl}(b)$ as the inclusion-smallest set in $S\cap\mathcal{P}(\mathrm{cl}(a))$ containing all points of $b$ and all limit points of $b$. ∎
@@
This is proved by transfinite induction on rank of $b$ in the inclusion ordering of $S$. If $b$ is a singleton, this is part of the initial assumptions on the posets $\mathbb{P}_{\alpha}$. Now, suppose that $b\in S$ is arbitrary and $b\subseteq N\cap\omega_{n}$, and the statement has been proved for all sets in $S$ of lower rank. Let $\bar{a}$ be a finite tuple of sets in $S$ such that $\bar{a}\subseteq N\cap\omega_{n}$. We will show that for every $\mathbf{Q}_{b,\bar{a}}$-name $\dot{C}$ for an open dense subset of $\omega^{<\omega}$ and $p\in\mathbf{Q}_{b,\bar{a}}$, there is a condition $q\leq p$ and a dense subset $D\subset\omega^{<\omega}$ such that $q\Vdash\check{D}\subset\dot{C}$.
@@
Let $\{c_{n}\colon n\in\omega\}$ be an enumeration of all proper subsets of $b$ which are in $S$. Define $\bar{a}_{0}=\bar{a}$ and for $n>0$, $\bar{a}_{n+1}=\bar{a}_{n}^{\frown}c_{n}$. Note that we can assume that for each $n\in\omega$, if $c_{n}\setminus\bigcup\bar{a}_{n}=\emptyset$, then $c_{n+1}\setminus\bigcup\bar{a}_{n+1}\neq\emptyset$. Now, for each $n\in\omega$ and $r\in\mathbf{Q}_{b,\bar{a}_{n}}$, let $\pi^{n}_{0}(r)=r\upharpoonright c_{n}$ and $\pi_{1}^{n}(r)=r\setminus\pi_{0}^{n}(r)$. Let $\{t_{n}\colon n\in\omega\}$ be an enumeration of $\omega^{<\omega}$. Now we build sequences $\langle\dot{C}_{n},D_{n},p_{n},q_{n},\rho_{n},s_{n}:n\in\omega\rangle$ as follows:
@@
    Let $q_{0}=p$, $\dot{C}_{0}=\dot{C}$, $D_{0}$ is the open dense set above defined, and $(\rho_{0},s_{0})\in D_{0}$ is such that $(\rho_{0},s_{0})\leq(q_{0},t_{0})$. Now define $p_{0}=\emptyset$.
@@
    For $n>0$ we have two cases:
    1. (a)
      $\pi_{0}^{n-1}[\mathbf{Q}_{b,\bar{a}_{n-1}}]=\emptyset$. Note that this happens if and only if $c_{n-1}\subseteq\bigcup\bar{a}_{n-1}$. In this case we define $D_{n}=D_{n-1}$ and $\dot{C}_{n}=\dot{C}_{n-1}$. Now extend the condition $(q_{n-1},t_{n})$ to find $(\rho_{n},s_{n})\in D_{n}$ such that $(\rho_{n},s_{n})\leq(q_{n-1},t_{n})$. Then define $q_{n}=\rho_{n}$ and $p_{n}=\emptyset$.
    2. (b)
      $\pi_{0}^{n-1}[\mathbf{Q}_{b,\bar{a}_{n-1}}]\neq\emptyset$. Assume everything is defined up to $n-1$. Let $q_{n}=\pi_{1}^{n-1}(\rho_{n-1})$. Define $\dot{C}_{n}$ to be the $\mathbf{Q}_{c_{n},\bar{a}_{n-1}}$-name as:
@@
For item (1) of the theorem, suppose that $q\in\mathbf{P}$ is a condition and $\tau$ is a $\mathbf{P}$-name such that $q\Vdash\tau\colon\omega\to\check{\omega}_{1}$ is a function. Use the correctness assumption on the true ground to find a countable elementary submodel $N$ of a large structure containing $q,\tau$ such that $N\cap S\in M$. Let $\langle O_{n}\colon n\in\omega\rangle$ enumerate open dense subsets of $\mathbf{P}$ in the model $N$. Let $\sigma$ be a $\mathbf{P}\cap N$-name defined by $\sigma=\{\langle r,\check{t}\rangle\colon$ for some number $n\in\omega$, $r\in O_{n}\cap N$ and $t$ is an integer sequence starting with $n\}$. Then $\mathbf{P}\cap N$ forces that $\sigma\subset\omega^{<\omega}$ is open dense. Apply Claim 8.20 with $\bar{a}=0$ and $b=N\cap\kappa$ to find a condition $p\leq q$ in $\mathbf{P}$ and an open dense subset $D\subset\omega^{<\omega}$ such that $p\Vdash_{\mathbf{P}}\check{D}\subset\sigma$. Then $p\in\mathbf{P}$ is an $N$-master condition; in particular, $p\Vdash\mathrm{rng}(\tau)\subset\check{N}\cap\omega_{1}$ and and (1) follows.
@@
(2) Follows from the fact that $\mathbb{P}_{\omega_{2}}(\dot{\mathcal{U}}_{\alpha_{\beta}})^{V[G_{\alpha_{\beta}}]}=\bigcup_{\gamma\in[\omega_{1},\omega_{2})}h_{\gamma}^{*}[\mathbb{P}_{\omega_{1}}(\dot{\mathcal{U}}_{\alpha_{\beta}})^{V[G_{\alpha_{\beta}}]}]$, that each condition in $\overline{\mathbf{P}}^{\omega_{2}}$ has countable domain, that $M_{b}$ satisfies $\mathsf{CH}$ for all $b\in S$. ∎
@@
$\bigcup_{j\leq n}E^{q(0)}_{j}\subseteq\bigcup_{l<k_{n}}E^{p(0)}_{l}$
note that such $k_{n}$ exists because $q$ is a simple condition. Now define a condition $r\in\mathbb{P}_{\omega_{2}}(\dot{\mathcal{U}}_{\alpha_{\beta}})^{V[G_{\alpha_{\beta}}]}$ as follows:
@@
(2) The proof is by induction on the rank of $b\in S$. If $b=\{\beta\}$, this is Corollary 8.25. Fix $b\in S$ and assume the assertion holds for all $d\in S$ having smaller rank than the rank of $b$. Let $\{c_{n}:n\in\omega\}$ be an enumeration of $\{c\in S:c\subseteq b\land c\neq b\}$. Define $\overline{a}_{0}=\{d\cap b:d\in\overline{a}\}$, and for $n>0$, $\overline{a}_{n}={\overline{a}_{n-1}}^{\frown}c_{n-1}$. Let $p\in\overline{\mathbf{P}}_{b,\overline{a}}^{\delta}$ be a condition stronger than $q\upharpoonright\upharpoonright\delta$. Now construct sequences $\langle p_{n},q_{n},r_{n}:n\in\omega\rangle$ as follows:
@@
    If $d_{n}=c_{n}\setminus\bigcup\overline{a}_{n}=\emptyset$. Define $r_{n}=\emptyset$, $p_{n}=p_{n-1}$ and $q_{n}=q_{n-1}$.
  4. (4)
    If $d_{n}=c_{n}\setminus\bigcup\overline{a}_{n}\neq\emptyset$. By Lemma 8.26, we have $q_{n-1}\upharpoonright c_{n}\in\overline{\mathbf{P}}^{\omega_{2}}_{c_{n},\overline{a}_{n}}$, $p_{n-1}\upharpoonright c_{n}\in\overline{\mathbf{P}}^{\delta}_{c_{n},\overline{a}_{n}}$, and $p_{n-1}\upharpoonright c_{n}\leq(q_{n-1}\upharpoonright c_{n})\upharpoonright\upharpoonright\delta$. By induction hypothesis we can find a condition $r\in\overline{\mathbf{P}}_{c_{n},\overline{a}_{n}}^{\omega_{2}}$ extending both $p_{n-1}\upharpoonright c_{n}$ and $q_{n-1}\upharpoonright c_{n}$; let $r_{n}$ be one of such conditions. Now define $p_{n}=p_{n-1}\upharpoonright(\omega_{2}\setminus c_{n})$ and $q_{n}=q_{n-1}\upharpoonright(\omega_{2}\setminus c_{n})$. Note that $p_{n}\in\overline{\mathbf{P}}_{b,\overline{a}_{n}^{\frown}c_{n}}^{\delta}$, $q_{n}\in\overline{\mathbf{P}}^{\omega_{2}}_{b,\overline{a}_{n}^{\frown}c_{n}}$ and $p_{n}\leq q_{n}\upharpoonright\upharpoonright\delta$.
Now define $p_{\omega}=\bigcup_{n\in\omega}r_{n}$. Let us see that $p_{\omega}$ is a condition in $\overline{\mathbf{P}}^{\omega_{2}}_{b,\overline{a}}$. Pick any $d\in S$. If $d\cap(b\setminus\bigcup\overline{a})=\emptyset$ then $p_{\omega}\upharpoonright d=\emptyset\in M_{d}$. Assume otherwise and let $n\in\omega$ be such that $c_{n}=d\cap(b\setminus\overline{a})$. This implies that $p_{\omega}\upharpoonright d=p_{\omega}\upharpoonright c_{n}=\bigcup_{j\leq n}r_{j}\upharpoonright c_{n}$. By induction hypothesis, each $r_{j}$ is a condition in $\overline{\mathbf{P}}_{c_{j},\overline{a}_{j}}^{\omega_{2}}\subseteq\overline{\mathbf{P}}^{\omega_{2}}$, which implies that $r_{j}\upharpoonright c_{n}\in M_{c_{n}}$; thus, $p_{\omega}\upharpoonright c_{n}=\bigcup_{j\leq n}r_{j}\upharpoonright c_{n}\in M_{c_{n}}$. Therefore, $p_{\omega}$ is a condition in $\overline{\mathbf{P}}^{\omega_{2}}$. Since $dom(p_{\omega})\subseteq b\setminus\bigcup\overline{a}$, we have that $p_{\omega}$ is a condition in $\overline{\mathbf{P}}^{\omega_{2}}_{b,\overline{a}}$. By construction we have that $p_{\omega}$ extends both $p$ and $q$.
@@
For any $\delta\in[\omega_{1},\omega_{2})$ and $b\in S$ such that $\beta\in S$, $M_{b}\vDash\overline{\mathbf{P}}^{\delta}_{b}\lessdot\overline{\mathbf{P}}^{\omega_{2}}_{b}$ and $V[G_{\omega_{2}}]\vDash\overline{\mathbf{P}}^{\delta}_{b}\lessdot\overline{\mathbf{P}}^{\omega_{2}}_{b}$.
@@
We proceed as in the $\omega_{1}$-case. Fix $\{p_{\alpha}:\alpha\in\omega_{2}\}\subseteq\overline{\mathbf{P}}^{\omega_{2}}$. We can assume that for each $\alpha$, $dom(p_{\alpha})\in S$. Let now $\langle N_{\alpha}:\alpha\in\omega_{1}\rangle$ be an $\in$-continuous sequence of countable elementary submodels such that $S,\omega_{1},\omega_{2},\overline{\mathbf{P}}^{\omega_{2}},\{p_{\alpha}:\alpha\in\omega_{2}\}\in N_{0}$, and for each $\alpha\in\omega_{1}$, $N_{\alpha}\cap\omega_{2}\in S$. Define $N_{\omega_{1}}=\bigcup_{\beta\in\omega_{1}}N_{\beta}$. First note that $\overline{\mathbf{P}}^{\omega_{2}}\cap N_{\omega_{1}}=\bigcup_{\gamma\in[\omega_{1},\omega_{2})\cap N_{\omega_{1}}}\overline{\mathbf{P}}^{\gamma}\cap N_{\omega_{1}}$.
Pick $\alpha_{0}\in\omega_{2}\setminus N_{\omega_{1}}$ and let $A=dom(p_{\alpha_{0}})\cap N_{\omega_{1}}$ and define $q=p_{\alpha_{0}}\upharpoonright A$. Note that $q\in\overline{\mathbf{P}}^{\omega_{2}}$ and $q\subseteq N_{\omega_{1}}$. Also, we can find $\gamma_{0}\in[\omega_{1},\omega_{2})\cap N_{\omega_{1}}$ such that for each $\delta\in A$, $q(\delta)\in\mathbb{P}_{\gamma_{0}}(\dot{\mathcal{U}}_{\alpha_{\delta}})^{V[G_{\alpha_{\delta}}]}$. For such $\gamma_{0}$, $\overline{\mathbf{P}}^{\gamma_{0}}\in N_{\omega_{1}}$ holds true, and $q\in\overline{\mathbf{P}}^{\gamma_{0}}_{A}$. However, $\overline{\mathbf{P}}^{\gamma_{0}}$ has cardinality $\omega_{1}$ and $N_{\omega_{1}}$ knows this, which implies $\overline{\mathbf{P}}^{\gamma_{0}}\subseteq N_{\omega_{1}}$, from which it follows that $q\in N_{\omega_{1}}$. By elementarity, we can find $\alpha_{1}\in\omega_{2}$ such that $p_{\alpha_{1}}\in N_{\omega_{1}}$ and $q\subseteq p_{\alpha_{1}}$. Therefore, $p_{\alpha_{0}}$ and $p_{\alpha_{1}}$ are compatible.

### v2

@@ 4. $\text{IX}_{n}$)
Let $x\subseteq\kappa$ be an infinite subset having order type $\omega$ and $f:x\to 2$. Since $\mathbb{Q}_{\kappa}$ is Cohen preserving, by Lemma 4.22 we find $\vec{\nu}\in\mathcal{C}_{\kappa}\cap V$ such that $\bigcap_{\alpha\in{x}}{B}^{\alpha}_{{f}(\alpha)}\subseteq{X}_{\vec{\nu}_{0}}$. By the last remark of the previous paragraph, ${\varphi}^{-1}_{\alpha_{\gamma}}[{X}_{\vec{\nu}_{0}}]\in{\mathcal{U}}_{\alpha_{\gamma}}^{*}$, which implies ${\varphi}^{-1}_{\alpha_{\gamma}}[\bigcap_{\alpha\in{x}}B^{\alpha}_{{f}(\alpha)}]\in{\mathcal{U}}_{\alpha_{\gamma}}^{*}$. Thus, $\bigcap_{\alpha\in{x}}{D}^{\alpha}_{{f}(\alpha)}\in{\mathcal{U}}_{\alpha_{\gamma}}^{*}$. Note that this implies that any ultrafilter extending $\mathcal{U}_{\alpha_{\gamma}}$ is Tukey top. Since $\mathcal{U}_{\alpha_{\gamma}}=\mathcal{F}\cap V[H_{\alpha_{\gamma}}]\subseteq\mathcal{F}\subseteq\mathcal{U}$, we have that $\mathcal{U}$ is Tukey top. ∎
@@ 10. Squares and products
In this section we present a different approach based on the existence of square sequences. Note that, working in the Sacks model, the quotient forcing $[\mathbb{Q}_{\kappa}:\mathbb{S}_{\kappa}]$ is naturally equivalent to a subposet of the countable support product $\bigotimes_{\beta\in S}\dot{\mathbb{P}}_{\kappa}(\dot{\mathcal{U}}_{\beta})$. In this section we provide a construction of a reduced product which keeps all the relevant properties of $\mathbb{Q}_{\kappa}$, although it looks quite different in what the support of the conditions refers. This construction is strongly based on Theorem 10.1 below, which is interesting by itself. The ingenious proof of Theorem 10.1 is due to the work of the second author. First we deal with the proof of this theorem and then we show how to apply it to construct a model where there all ultrafilters are Tukey top.
@@
Suppose that $n\in\omega$ is a natural number and $\Box_{m}$ holds for all $m<n$ and $\mathtt{cof}(meager)=\aleph_{1}$ holds. Let $\langle\mathbb{P}_{\alpha}\colon\alpha\in\omega_{n}\rangle$ be a sequence of $\sigma$-proper, Cohen preserving posets of cardinality $\aleph_{1}$. Then there is a poset $\mathbf{P}$ which:
@@
If $V=L[\vec{B}]$ and $\vec{B}$ is an $\omega_{1}$-sequence of ordinals, then $\mathsf{GCH}$ holds.
@@
The model $M_{a}$ is $L[K][S\cap\mathcal{P}(\mathrm{cl}(a))][\langle\langle\alpha,\mathbb{P}_{\alpha}\rangle\colon\alpha\in a\rangle]$. One point here is that this is a model of ZFC, since the set $S\cap\mathcal{P}(\mathrm{cl}(a))$ is injectively mapped into $Ord^{n+1}$ by the function $s$ and thereby canonically well-ordered. Items (1, 2) are obvious. For item (3), note that $S\cap\mathcal{P}(cl(a))$ is a countable set, so $M_{a}=L[\vec{B}_{a}]$ for some $\omega_{1}$-sequence $\vec{B}_{a}$ of ordinals and it computes $\aleph_{1}$ correctly by the choice of the sequence $K$. It follows from 10.10 that $M_{a}$ is a model of $\mathsf{CH}$. In the model $M_{a}$, let $\prec_{a}$ be the first well-ordering of the reals of ordertype $\omega_{1}$ in the natural constructibility well-order. For the last item, note that if $b\subseteq a$ are sets in $S$, then $M_{a}$ can compute $\mathrm{cl}(b)$ as the inclusion-smallest set in $S\cap\mathcal{P}(\mathrm{cl}(a))$ containing all points of $b$ and all limit points of $b$. ∎
@@
This is proved by transfinite induction on rank of $b$ in the inclusion ordering of $S$. If $b$ is a singleton, this is part of the initial assumptions on the posets $\mathbb{P}_{\alpha}$. Now, suppose that $b\in S$ is arbitrary and $b\subseteq N\cap\omega_{n}$, and the statement has been proved for all sets in $S$ of lower rank. Let $\bar{a}$ be a finite tuple of sets in $S$ such that $\bar{a}\subseteq N\cap\omega_{n}$. We will show that for every $\mathbf{Q}_{b,\bar{a}}$-name $\dot{C}$ for an open dense subset of $\omega^{<\omega}$ and $q\in\mathbf{Q}_{b,\bar{a}}$, there is a condition $p\leq q$ and a dense subset $D\subset\omega^{<\omega}$ such that $p\Vdash\check{D}\subset\dot{C}$.
@@
Now we work in $M_{b}$. Let $\{c_{n}\colon n\in\omega\}$ be an enumeration of all proper subsets of $b$ which are in $S$. Define $\bar{a}_{0}=\bar{a}$ and for $n>0$, $\bar{a}_{n+1}=\bar{a}_{n}^{\frown}c_{n}$. By going to a subsequece if necessary, we can assume that:
@@
    For each $n\in\omega$, $c_{n}\setminus\bigcup\bar{a}_{n}\neq\emptyset$.
@@
    For each $c\in S$ subset of $b$ such that $c\setminus\bigcup\bar{a}\neq\emptyset$, if $c$ does not appear in $\{c_{n}:n\in\omega\}$ then there is $k\in\omega$ such that $c\subseteq\bigcup_{l\leq k}c_{l}\cup\bigcup\bar{a}$.
Now, for each $n\in\omega$ and $r\in\mathbf{Q}_{b,\bar{a}_{n}}$, let $\pi^{n}_{0}(r)=r\upharpoonright c_{n}$ and $\pi_{1}^{n}(r)=r\setminus\pi_{0}^{n}(r)$, so
$\pi_{n}:\mathbf{Q}_{b,\bar{a}_{n}}\to\mathbf{Q}_{c_{n},\bar{a}_{n}}\times\mathbf{Q}_{b,\bar{a}_{n+1}}$
given by $\pi_{n}(p)=(\pi_{0}^{n}(p),\pi_{1}^{n}(p))$ is an order isomorphism. Let $\{t_{n}\colon n\in\omega\}$ be an enumeration of $\omega^{<\omega}$. Now we build sequences $\langle\dot{C}_{n},D_{n},p_{n},q_{n},\rho_{n},s_{n}:n\in\omega\rangle$ as follows:
  1. (1)
    Let $q_{0}=q$, $\dot{C}_{0}=\dot{C}$, $D_{0}$ is the open dense $E(\dot{C})$ set above defined, and $(\rho_{0},s_{0})\in D_{0}$ is such that $(\rho_{0},s_{0})\leq(q_{0},t_{0})$. Now define $p_{0}=\emptyset$.
  2. (2)
    For $n>0$, assume everything is defined up to $n-1$. Let $q_{n}=\pi_{1}^{n-1}(\rho_{n-1})$. Define $\dot{C}_{n}$ to be the $\mathbf{Q}_{c_{n},\bar{a}_{n-1}}$-name,
@@
For item (1) of the theorem, suppose that $q\in\mathbf{P}$ is a condition and $\dot{\tau}$ is a $\mathbf{P}$-name such that $q\Vdash\dot{\tau}\colon\omega\to\check{\omega}_{1}$ is a surjective function. Since In $V$ there are $\omega_{1}$ open dense subsets of $\omega^{<\omega}$ whose complements generate the ideal $\mathsf{nwd}$, $q$ should force the existence of an open dense set which diagonalizes each dense set from $V$. Let $\dot{D}$ be a $\mathbf{P}$-name for such set. In particular, we have that for all open dense set $E\subseteq\omega^{<\omega}$ in $V$, $q\Vdash E\setminus\dot{D}$ is infinite. Let $N$ be a countable elementary submodel such that $\mathbf{P},q,\dot{D}\in N$ and $b=N\cap\omega_{n}\in S$. Then, as before, we can find $p\leq q$ and an open dense set $D_{0}$ in $V$ such that $p\Vdash D_{0}\subseteq\dot{D}$, which is a contradiction.
@@
(2) Follows from the fact that $\mathbb{P}_{\omega_{2}}(\dot{\mathcal{U}}_{\alpha_{\beta}})^{V[G_{\alpha_{\beta}}]}=\bigcup_{\gamma\in[\omega_{1},\omega_{2})}h_{\gamma}^{*}[\mathbb{P}_{\omega_{1}}(\dot{\mathcal{U}}_{\alpha_{\beta}})^{V[G_{\alpha_{\beta}}]}]$, that each condition in $\overline{\mathbf{P}}^{\omega_{2}}$ has countable domain, and that $M_{b}$ satisfies $\mathsf{CH}$ for all $b\in S$. ∎
@@
$\bigcup_{j\leq n}E^{q(0)}_{j}\subseteq dom(h_{p(0)})\cup\bigcup_{l<k_{n}}E^{p(0)}_{l}$
note that such $k_{n}$ exists because $q$ is a simple condition. Note that for each $n\in\omega$, $\vec{E}_{p(0)}*k_{n}\sqsubseteq\vec{E}_{q(0)}*(n+1)=\vec{E}_{q(\gamma_{n})}$. Now define a condition $r\in\mathbb{P}_{\omega_{2}}(\dot{\mathcal{U}}_{\alpha_{\beta}})^{V[G_{\alpha_{\beta}}]}$ as follows:
@@
(2) The proof is by induction on the rank of $b\in S$. If $b=\{\beta\}$, this is Corollary 10.25. Fix $b\in S$ and assume the assertion holds for all $d\in S$ having smaller rank than the rank of $b$. We work in $M_{b}$. Let $\{c_{n}:n\in\omega\}$ be an enumeration of $\{c\in S:c\subseteq b\land c\neq b\}$. Define $\overline{a}_{0}=\{d\cap b:d\in\overline{a}\}$, and for $n>0$, $\overline{a}_{n}={\overline{a}_{n-1}}^{\frown}c_{n-1}$. By going to a subsequece if necessary, we can assume that:
  1. (1)
    For each $n\in\omega$, $c_{n}\setminus\bigcup\bar{a}_{n}\neq\emptyset$.
  2. (2)
    For each $c\in S$ subset of $b$ such that $c\setminus\bigcup\bar{a}\neq\emptyset$, if $c$ does not appear in $\{c_{n}:n\in\omega\}$ then there is $k\in\omega$ such that $c\subseteq\bigcup_{l\leq k}c_{l}\cup\bigcup\bar{a}$.
Let $p\in\overline{\mathbf{P}}_{b,\overline{a}}^{\delta}$ be a condition stronger than $q\upharpoonright\upharpoonright\delta$. Now construct sequences $\langle p_{n},q_{n},r_{n}:n\in\omega\rangle$ as follows:
@@
    By Lemma 10.26, we have $q_{n-1}\upharpoonright c_{n}\in\overline{\mathbf{P}}^{\omega_{2}}_{c_{n},\overline{a}_{n}}$, $p_{n-1}\upharpoonright c_{n}\in\overline{\mathbf{P}}^{\delta}_{c_{n},\overline{a}_{n}}$, and $p_{n-1}\upharpoonright c_{n}\leq(q_{n-1}\upharpoonright c_{n})\upharpoonright\upharpoonright\delta$. By induction hypothesis we can find a condition $r\in\overline{\mathbf{P}}_{c_{n},\overline{a}_{n}}^{\omega_{2}}$ extending both $p_{n-1}\upharpoonright c_{n}$ and $q_{n-1}\upharpoonright c_{n}$; let $r_{n}$ be one of such conditions. Now define $p_{n}=p_{n-1}\upharpoonright(\omega_{2}\setminus c_{n})$ and $q_{n}=q_{n-1}\upharpoonright(\omega_{2}\setminus c_{n})$. Note that $p_{n}\in\overline{\mathbf{P}}_{b,\overline{a}_{n}^{\frown}c_{n}}^{\delta}$, $q_{n}\in\overline{\mathbf{P}}^{\omega_{2}}_{b,\overline{a}_{n}^{\frown}c_{n}}$ and $p_{n}\leq q_{n}\upharpoonright\upharpoonright\delta$.
Now define $p_{\omega}=\bigcup_{n\in\omega}r_{n}$. Let us see that $p_{\omega}$ is a condition in $\overline{\mathbf{P}}^{\omega_{2}}_{b,\overline{a}}$. Pick any $d\in S$. If $d\cap(b\setminus\bigcup\overline{a})=\emptyset$ then $p_{\omega}\upharpoonright d=\emptyset\in M_{d}$. Assume otherwise and let $c=d\cap b$. By definition of the enumeration $\{c_{n}:n\in\omega\}$, there is $n\in\omega$ such that $c\subseteq\bigcup_{l\leq n}c_{l}\cup\bigcup\overline{a}$, so we have that $p_{\omega}\upharpoonright d=p_{\omega}\upharpoonright c=\bigcup_{j\leq n}r_{j}\upharpoonright c$. By induction hypothesis, each $r_{j}$ is a condition in $\overline{\mathbf{P}}_{c_{j},\overline{a}_{j}}^{\omega_{2}}\subseteq\overline{\mathbf{P}}^{\omega_{2}}$, which implies that $r_{j}\upharpoonright c\in M_{c}$; thus, $p_{\omega}\upharpoonright c=\bigcup_{j\leq n}r_{j}\upharpoonright c\in M_{c}$. Therefore, $p_{\omega}$ is a condition in $\overline{\mathbf{P}}^{\omega_{2}}$. Since $dom(p_{\omega})\subseteq b\setminus\bigcup\overline{a}$, we have that $p_{\omega}$ is a condition in $\overline{\mathbf{P}}^{\omega_{2}}_{b,\overline{a}}$. By construction we have that $p_{\omega}$ extends both $p$ and $q$.
@@
For any $\delta\in[\omega_{1},\omega_{2})$ and $b\in S$, $M_{b}\vDash\overline{\mathbf{P}}^{\delta}_{b}\lessdot\overline{\mathbf{P}}^{\omega_{2}}_{b}$ and $V[G_{\omega_{2}}]\vDash\overline{\mathbf{P}}^{\delta}_{b}\lessdot\overline{\mathbf{P}}^{\omega_{2}}_{b}$.
@@
We proceed as in the $\omega_{1}$-case. Fix $\{p_{\alpha}:\alpha\in\omega_{2}\}\subseteq\overline{\mathbf{P}}^{\omega_{2}}$. We can assume that for each $\alpha$, $dom(p_{\alpha})\in S$. Let now $\langle N_{\alpha}:\alpha\in\omega_{1}\rangle$ be an $\in$-increasing sequence of countable elementary submodels such that $S,\omega_{1},\omega_{2},\overline{\mathbf{P}}^{\omega_{2}},\{p_{\alpha}:\alpha\in\omega_{2}\}\in N_{0}$, and for each $\alpha\in\omega_{1}$, $N_{\alpha}\cap\omega_{2}\in S$. Define $N_{\omega_{1}}=\bigcup_{\beta\in\omega_{1}}N_{\beta}$. First note that $\overline{\mathbf{P}}^{\omega_{2}}\cap N_{\omega_{1}}=\bigcup_{\gamma\in[\omega_{1},\omega_{2})\cap N_{\omega_{1}}}\overline{\mathbf{P}}^{\gamma}\cap N_{\omega_{1}}$.
Pick $\alpha_{0}\in\omega_{2}\setminus N_{\omega_{1}}$, let $A=dom(p_{\alpha_{0}})\cap N_{\omega_{1}}$ and define $q=p_{\alpha_{0}}\upharpoonright A$. Since $A$ is countable, there is $\gamma\in\omega_{1}$ such that $A\subseteq N_{\gamma}$. Since $N_{\gamma}\cap\omega_{2}\in S$, we have that $A=dom(q_{\alpha_{0}})\cap B$, where $B=\omega_{2}\cap N_{\gamma}$. Moreover, since $S,N_{\gamma}\in N_{\gamma+1}$, we have $B\in N_{\gamma+1}$. Note that $q\in\overline{\mathbf{P}}^{\omega_{2}}$ and $q\subseteq N_{\omega_{1}}$ both hold true. Also, we can find $\gamma_{0}\in[\omega_{1},\omega_{2})\cap N_{\omega_{1}}$ such that for each $\delta\in A$, $q(\delta)\in\mathbb{P}_{\gamma_{0}}(\dot{\mathcal{U}}_{\alpha_{\delta}})^{V[G_{\alpha_{\delta}}]}$. For such $\gamma_{0}$, $\overline{\mathbf{P}}^{\gamma_{0}}\in N_{\omega_{1}}$ holds true, and we also have that $q\in\overline{\mathbf{P}}^{\gamma_{0}}_{B}\in N_{\omega_{1}}$. However, $\overline{\mathbf{P}}^{\gamma_{0}}_{B}$ has cardinality $\omega_{1}$ and $N_{\omega_{1}}$ knows this, which implies $\overline{\mathbf{P}}^{\gamma_{0}}_{B}\subseteq N_{\omega_{1}}$, from which it follows that $q\in N_{\omega_{1}}$. By elementarity, we can find $\alpha_{1}\in\omega_{2}$ such that $p_{\alpha_{1}}\in N_{\omega_{1}}$ and $q\subseteq p_{\alpha_{1}}$. Therefore, $p_{\alpha_{0}}$ and $p_{\alpha_{1}}$ are compatible.